OpenStudy (anonymous):

27^{5 x} = 9^{x+3}. Find X

OpenStudy (anonymous):

\(27^{5x} = 9^{x+3}\) It seems like you should make the bases the same, then set the exponents equal to each other and solve for x.

OpenStudy (anonymous):

any idea how i could make the bases the same. i tried but failed to do so.

OpenStudy (anonymous):

\(3^3=27\) and \(3^2=9\)

OpenStudy (anonymous):

Thanks alot

OpenStudy (anonymous):

You're Welcome. One last thought... \((3^3)^{5x}=3^{15x}\) When raising a power to a power, you multiply.

OpenStudy (anonymous):

Got it. Thank you again for your help